A fair coin is flipped 5 times. What is the probability of getting at least 4 tails?
Answer by Boreal(15235)

At least 4 tails = probability of 4 tails and probability of 5 tails.
For 4, there are 5 different ways(1/2)^4(1/2)=5/32
For 5 tails, it is (1/2)^5=1/32
Total is 6/32 or 3/16
Pascal's triangle for 5 flips is 1==5==10==10=5==1